OpenStudy (anonymous):

How do I go about solving this complex compound interest problem with a variable?

OpenStudy (anonymous):

\[50000=20000(1+\frac{ x }{ 12})^{25}\]

jimthompson5910 (jim_thompson5910):

The first step is to divide both sides by 20,000

jimthompson5910 (jim_thompson5910):

Doing that and then flipping the two sides gives you \[\Large \left(1+\frac{ x }{ 12}\right)^{25} = 2.5\]

jimthompson5910 (jim_thompson5910):

The next step is to raise both sides to the power 1/25 this will cancel out the exponent of 25 on the left side of my equation above (since 25*(1/25) = 1) So we'll have \[\Large 1+\frac{ x }{ 12} = (2.5)^{1/25}\]

jimthompson5910 (jim_thompson5910):

do you see how to finish up?

OpenStudy (anonymous):

yeah, that makes a ton of sense :O!!!! Thanks!!!

OpenStudy (anonymous):

Think you can help with one more :'O?

jimthompson5910 (jim_thompson5910):

sure go ahead

OpenStudy (anonymous):

\[50000=20000(1+\frac{ 0.0315 }{ 12})^{12*x}\]

jimthompson5910 (jim_thompson5910):

what is your first step

OpenStudy (anonymous):

Divide by 20000 so I end up with...\[2.5=(1+\frac{ 0.0315 }{ 12 })^{12*x}\]

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

now the next step is to somehow isolate that exponent 12x we use logs to do this

jimthompson5910 (jim_thompson5910):

apply the log base 10 to both sides \[\Large \log(2.5)=\log\left((1+\frac{ 0.0315 }{ 12 })^{12*x}\right)\] doing this allows us to pull down the exponent 12x \[\Large \log(2.5)=12x*\log\left(1+\frac{ 0.0315 }{ 12 }\right)\] what's next?

OpenStudy (anonymous):

\[0.4=12x+0.0114\]

OpenStudy (anonymous):

Right...?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

1+0.0315/12 = 1.002625 the log of that is log(1.002625) = 0.00113852934813 which is roughly 0.00114 which is what you got

jimthompson5910 (jim_thompson5910):

however, it should be 12x*0.00114 on the right side

jimthompson5910 (jim_thompson5910):

also, on the left side, I would use at least 3 decimal digits of accuracy

jimthompson5910 (jim_thompson5910):

what I got so far is 0.3979 = 12x*0.0011385 do you see how to finish up?

OpenStudy (anonymous):

x=29.12458?

OpenStudy (anonymous):

About?

jimthompson5910 (jim_thompson5910):

yes approximately

OpenStudy (anonymous):

Ah ok! That makes a lot of sense!

OpenStudy (anonymous):

I really can't thank you enough, you really explained it in a very understandable way.

jimthompson5910 (jim_thompson5910):

I'm glad things are clicking now